OpenStudy (anonymous):

Two cars leave towns 750 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour more than the other's. If they meet in 5 hours, what is the rate of the faster car? Do not do any rounding.

OpenStudy (anonymous):

18x*5x=750

ganeshie8 (ganeshie8):

this is slightly different from previous problem we need to be a lil bit careful here

ganeshie8 (ganeshie8):

see you're not given speeds of cars, you're given how much faster one car is compared to the other

ganeshie8 (ganeshie8):

say the speed of faster car is \(x\) then the speed of slower car would be \(x-18\) yes ?

OpenStudy (anonymous):

18 and 5

OpenStudy (anonymous):

yes?

ganeshie8 (ganeshie8):

and you know they travel exactly 750 kilometers by the time they meet (in 5 hours)

ganeshie8 (ganeshie8):

\[5x + 5(x-18) = 750\] solve \(x\)

OpenStudy (anonymous):

5x+5x-90=750 10x-90=750 10x=840 x=84

ganeshie8 (ganeshie8):

yes thats the speed of faster car

OpenStudy (anonymous):

thanks i think thats it for know u are great talk to u later

ganeshie8 (ganeshie8):

you're welcome ! have good sleep :)